# Chain-of-Factors Paper-Reviewer Matching

### v2

@@ Abstract.
With the rapid increase in paper submissions to academic conferences, the need for automated and accurate paper-reviewer matching is more critical than ever. Previous efforts in this area have considered various factors to assess the relevance of a reviewer’s expertise to a paper, such as the semantic similarity, shared topics, and citation connections between the paper and the reviewer’s previous works. However, most of these studies focus on only one factor, resulting in an incomplete evaluation of the paper-reviewer relevance. To address this issue, we propose a unified model for paper-reviewer matching that jointly considers semantic, topic, and citation factors. To be specific, during training, we instruction-tune a contextualized language model shared across all factors to capture their commonalities and characteristics; during inference, we chain the three factors to enable step-by-step, coarse-to-fine search for qualified reviewers given a submission. Experiments on four datasets (one of which is newly contributed by us) spanning various fields such as machine learning, computer vision, information retrieval, and data mining consistently demonstrate the effectiveness of our proposed Chain-of-Factors model in comparison with state-of-the-art paper-reviewer matching methods and scientific pre-trained language models.111 Code, datasets, and our pre-trained model are available upon request now and will be published online later.
@@
ccs: Information systems Retrieval models and rankingccs: Computing methodologies Natural language processing
@@ 1. Introduction
Automatic paper-reviewer matching has become an increasingly crucial task recently due to explosive growth in the number of submissions to conferences and journals. Given a huge volume of (e.g., several thousand) submissions, it is prohibitively time-consuming for chairs or editors to manually assign papers to appropriate reviewers. Even if reviewers can self-report their expertise on certain papers through a bidding process, they can hardly scan all submissions, hence an accurate pre-ranking result should be delivered to them so that they just need to check a shortlist of papers. In other words, a precise scoring system that can automatically judge the expertise relevance between each paper and each reviewer becomes an increasingly urgent need for finding qualified reviewers.
@@
Paper-reviewer matching has been extensively studied as a text mining task (Mimno and McCallum, 2007; Charlin and Zemel, 2013; Jin et al., 2017; Anjum et al., 2019; Singh et al., 2023; Stelmakh et al., 2023), which aims to estimate to what extent a reviewer is qualified to review a submission given the text (e.g., title and abstract) of the submission as well as the papers previously written by the reviewer. Intuitively, as shown in Figure 1, there are three major factors considered by related studies. (1) Semantic: Taking the submission $p$ as a query, if the papers most semantically relevant to the query are written by a reviewer $r$, then $r$ should be qualified to review $p$. This intuition is used by previous methods such as the Toronto Paper Matching System (TPMS) (Charlin and Zemel, 2013), where tf–idf is used for calculating the semantic relevance. (2) Topic: If a reviewer $r$’s previous papers share many fine-grained research topics with the submission $p$, then $r$ is assumed to be an expert reviewer of $p$. This assumption is utilized by topic modeling approaches (Mimno and McCallum, 2007; Jin et al., 2017; Anjum et al., 2019). (3) Citation: Authors of the papers cited by the submission $p$ are more likely to be expert reviewers of $p$. This intuition is leveraged by studies (Singh et al., 2023; Stelmakh et al., 2023) using citation-enhanced scientific pre-trained language models (PLMs) such as SPECTER (Cohan et al., 2020) and SciNCL (Ostendorff et al., 2022). Note that the majority of previous studies do not assume that the topics and references of each paper are provided as input. Instead, such information should be inferred from the paper text. [^2] The reasons why related studies make such an assumption are multifold in our view. To be specific, topics selected by the authors when they submit the paper are too coarse (e.g., “Text Mining”), while paper-reviewer matching relies heavily on more fine-grained topics (e.g., “Community Question Answering”); references in the submission do not necessarily cover all papers that ought to be cited, so we should infer what the submission should cite rather than what it actually cites.
@@
We conduct experiments on four datasets covering different fields including machine learning, computer vision, information retrieval, and data mining. Three of the datasets are released in previous studies (Mimno and McCallum, 2007; Singh et al., 2023; Karimzadehgan et al., 2008). The fourth is newly annotated by us, which is larger than the previous three and contains more recent papers. Experimental results show that our proposed CoF model consistently outperforms state-of-the-art paper-reviewer matching approaches and scientific PLM baselines on all four datasets. Further ablation studies validate the reasons why CoF is effective: (1) CoF jointly considers three factors rather than just one, (2) CoF chains three factors to facilitate step-by-step, coarse-to-fine relevant paper selection instead of merging all factors in one step, and (3) CoF improves upon the baselines in each factor empirically.
@@ 2.1. Problem Definition
Given a set of paper submissions $\mathcal{P}=\{p_{1},p_{2},...,p_{M}\}$ and a set of candidate reviewers =$\mathcal{R}=\{r_{1},r_{2},...,r_{N}\}$, the paper-reviewer match- ing task aims to learn a function $f:\mathcal{P}\times\mathcal{R}\rightarrow\mathbb{R}$, where $f(p,r)$ reflects the expertise relevance between the paper $p$ and the reviewer $r$ (i.e., how knowledgeable that $r$ is to review $p$). We conform to the following three key assumptions made by previous studies (Mimno and McCallum, 2007; Karimzadehgan et al., 2008; Liu et al., 2014; Anjum et al., 2019; Singh et al., 2023): (1) We do not know any $f(p,r)\ (p\in\mathcal{P},r\in\mathcal{R})$ as supervision, which is a natural assumption for a fully automated paper-reviewer matching system. In other words, $f$ should be derived in a zero-shot setting, possibly by learning from available data for other tasks. (2) To characterize each paper $p\in\mathcal{P}$, its text information (e.g., title and abstract) is available, denoted by $\textsf{Text}(p)$. (3) To characterize each reviewer $r\in\mathcal{R}$, its previous papers are given, denoted by $\mathcal{Q}_{r}=\{q_{r,1},q_{r,2},...,q_{r,|\mathcal{Q}_{r}|}\}$. The text information of each previous paper $q\in\mathcal{Q}_{r}$ is also provided. $\mathcal{Q}_{r}$ is called the publication profile of $r$ (Mysore et al., 2023). In practice, $\mathcal{Q}_{r}$ may be a subset of $r$’s previous papers (e.g., those published within the last 10 years or those published in top-tier venues only). To summarize, the task is defined as follows:
@@ 2.2. Semantic, Topic, and Citation Factors
Citation Factor. Recent studies (Singh et al., 2023; Stelmakh et al., 2023) adopt scientific PLMs, such as SPECTER (Cohan et al., 2020) and SciNCL (Ostendorff et al., 2022), for paper-reviewer matching. During their pre-training, both SPECTER and SciNCL are initialized from SciBERT (Beltagy et al., 2019) and trained on a large number of citation links between papers. Empirical results show that emphasizing such citation information significantly boosts their performance in comparison with SciBERT. The motivation of considering the citation factor in paper-reviewer matching is also clear: if a paper $p$ cites many papers written by a reviewer $r$, then $r$ is more likely to be a qualified reviewer of $p$.
@@ 3.1. Chain-of-Factors Matching
We then rank all papers in $\mathcal{Q}_{\mathbb{S}}$ according to $f_{\rm topic}(p,\cdot)$ and pick those top-ranked ones as the output of this step, which we denote as $\mathcal{Q}_{\mathbb{S}\rightarrow\mathbb{T}}$, where $\mathbb{S}$ stands for the semantic factor.
After checking the topic factor, we further consider citation signals. Given that two papers share common fine-grained research topics, the citation link should provide an even stronger signal of the relevance between two papers. For instance, if two papers are both about “Information Extraction”, then one citing the other may further imply that they are studying the same task or using the same dataset. However, without the premise that two papers have common research fields, the citation link becomes a weaker indicator. For example, a paper about “Information Extraction” can cite a paper about “Large Language Models” simply because the former paper uses the large language model released in the latter one. This highlights our motivation to chain the three factors for step-by-step, coarse-to-fine selection of relevant papers and expert reviewers. Formally, given $\mathcal{Q}_{\mathbb{S}\rightarrow\mathbb{T}}$, we use a citation-related instruction to calculate the citation-aware relevance between $p$ and each selected paper.
@@
Here, $f(p,r)$ is the final relevance score between $p$ and $r$, which can be used to rank all candidate reviewers for $p$. Note that in the last step, we consider the sum of three types of relevance, so our chain-of-factors matching strategy can be denoted as $\mathbb{S}\rightarrow\mathbb{T}\rightarrow\mathbb{S}+\mathbb{T}+\mathbb{C}$. In our experiments (Section 4.3), we will demonstrate its advantage over only considering the citation factor in the last step (i.e., $\mathbb{S}\rightarrow\mathbb{T}\rightarrow\mathbb{C}$).
@@ 3.3. Model Training
In this section, we introduce the data and objective used to pre-trained the instruction-guided paper encoder $g(\cdot|\phi)$.
Pre-training Data. For the semantic factor, each submission $p$ is treated as a “query” and each paper $q$ in a reviewer’s publication profile is viewed as a “document”. In this case, $g(\cdot|\phi)$ should be learned to maximize the inner product of an ad-hoc query and its semantically relevant document in the vector space. To facilitate this, we adopt the Search dataset from the SciRepEval benchmark (Singh et al., 2023) to pre-train our model, where the queries are collected from an academic search engine, and the relevant documents are derived from large-scale user click-through data.
For the topic factor, $g(\cdot|\phi)$ should be trained to maximize the inner product of two papers $p$ and $q$ if they share common research fields. We utilize the MAPLE benchmark (Zhang et al., 2023b) as pre-training data in which millions of scientific papers are tagged with their fine-grained fields-of-study from the Microsoft Academic Graph (Sinha et al., 2015). For example, for CS papers in MAPLE, there are over 15K fine-grained research fields, and each paper is tagged with about 6 fields on average. Such data are used to derive topically relevant paper pairs to pre-train $g(\cdot|\phi)$.
For the citation factor, $g(\cdot|\phi)$ should be learned to maximize the inner product of two papers $p$ and $q$ if $p$ cites $q$. Following (Cohan et al., 2020; Ostendorff et al., 2022), we leverage a large collection of citation triplets $(p,q^{+},q^{-})$ constructed by Cohan et al. (Cohan et al., 2020), where $p$ cites $q^{+}$ but does not cite $q^{-}$, to pre-train $g(\cdot|\phi)$.
@@
Pre-training Objective. For all three factors, each sample from their pre-training data can be denoted as $(p,q^{+},q^{-}_{1},q^{-}_{2},...,q^{-}_{T})$, where $q^{+}$ is relevant to $p$ (i.e., when $\phi={\rm semantic}$, $q^{+}$ is clicked by users in a search engine given the search query $p$; when $\phi={\rm topic}$, $q^{+}$ shares fine-grained research fields with $p$; when $\phi={\rm citation}$, $q^{+}$ is cited by $p$) and $q^{-}_{t}\ (t=1,2,...,T)$ are irrelevant to $p$. Given a factor $\phi$ and its training sample $(p,q^{+},q^{-}_{1},q^{-}_{2},...,q^{-}_{T})$, we randomly pick an instruction $i_{\phi}$ corresponding to $\phi$. Then, using the instruction encoder ${\rm Enc}_{i}(\cdot)$ and the paper encoder ${\rm Enc}_{p}(\cdot)$, we can obtain $g(p|\phi)$, $g(q^{+}|\phi)$, $g(q^{-}_{1}|\phi)$, …, and $g(q^{-}_{T}|\phi)$, and we adopt a contrastive loss (Oord et al., 2018) to train our model:
@@
Hard Negatives. Cohan et al. (Cohan et al., 2020) show that a combination of easy negatives and hard negatives boosts the performance of their contrastive learning model. Following their idea, we take in-batch negatives (Karpukhin et al., 2020) as easy negatives and adopt the following strategies to find hard negatives: when $\phi={\rm semantic}$, $q^{-}_{t}$ is a hard negative if it is shown to users but not clicked given the query $p$; when $\phi={\rm topic}$, $q^{-}_{t}$ is a hard negative if it shares the same venue but does not share any fine-grained field with $p$; when $\phi={\rm citation}$, $q^{-}_{t}$ is a hard negative if $q^{+}$ cites $q^{-}_{t}$ but $p$ does not cite $q^{-}_{t}$.
@@ 4.1.1. Evaluation Datasets
  • NIPS (Mimno and McCallum, 2007) is a pioneering benchmark dataset for evaluating paper-reviewer matching. It consists of expertise relevance judgements between 34 papers accepted by NIPS 2006 and 190 reviewers. Annotations were done by 9 researchers from the NIPS community, and the score of each annotated paper-reviewer pair can be “3” (very relevant), “2” (relevant), “1” (slightly relevant), or “0” (irrelevant). Note that for each paper, the annotators only judge its relevance with a subset of reviewers, and a total of 393 paper-reviewer pairs are annotated.
  • SciRepEval (Singh et al., 2023) is a comprehensive benchmark for scientific document representation learning. Its paper-reviewer matching dataset combines the annotation effort from multiple sources (Mimno and McCallum, 2007; Liu et al., 2014; Zhao et al., 2022). Specifically, Liu et al. (Liu et al., 2014) added relevance scores of 766 paper-reviewer pairs to the NIPS dataset to mitigate its annotation sparsity; Zhao et al. (Zhao et al., 2022) provided 694 paper-reviewer relevance ratings for the ICIP 2016 conference. The combined dataset still adopts the “0”-“3” rating scale.
  • SIGIR (Karimzadehgan et al., 2008) contains 73 papers accepted by SIGIR 2007 and 189 prospective reviewers. Instead of annotating each specific paper-reviewer pair, the dataset constructors assign one or more aspects of information retrieval (e.g., “Evaluation”, “Web IR”, and “Language Models”, with 25 candidate aspects in total) to each paper and each reviewer. Then, the relevance between a paper and a reviewer is determined by their aspect-level similarity. In our experiments, to align with the rating scale in NIPS and SciRepEval, we discretize the Jaccard similarity between a paper’s aspects and a reviewer’s aspects to map their relevance to “0”, “1”, “2”, or “3”.

### v3

@@ Abstract.
With the rapid increase in paper submissions to academic conferences, the need for automated and accurate paper-reviewer matching is more critical than ever. Previous efforts in this area have considered various factors to assess the relevance of a reviewer’s expertise to a paper, such as the semantic similarity, shared topics, and citation connections between the paper and the reviewer’s previous works. However, most of these studies focus on only one factor, resulting in an incomplete evaluation of the paper-reviewer relevance. To address this issue, we propose a unified model for paper-reviewer matching that jointly considers semantic, topic, and citation factors. To be specific, during training, we instruction-tune a contextualized language model shared across all factors to capture their commonalities and characteristics; during inference, we chain the three factors to enable step-by-step, coarse-to-fine search for qualified reviewers given a submission. Experiments on four datasets (one of which is newly contributed by us) spanning various fields such as machine learning, computer vision, information retrieval, and data mining consistently demonstrate the effectiveness of our proposed Chain-of-Factors model in comparison with state-of-the-art paper-reviewer matching methods and scientific pre-trained language models. Code and Datasets are available at https://github.com/yuzhimanhua/CoF.
@@
ccs: Information systems Retrieval models and rankingccs: Computing methodologies Natural language processingjournalyear: 2025copyright: ccconference: Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Proceedings of the ACM Web Conference 2025 (WWW ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3696410.3714708isbn: 979-8-4007-1274-6/25/04
[…]
by
@@ 1. Introduction
Finding experts with certain knowledge in online communities has wide applications on the Web, such as community question answering on Stack Overflow and Quora (Riahi et al., 2012; Fu et al., 2020) as well as authoritative scholar mining from DBLP and AMiner (Deng et al., 2008; Zhang et al., 2008). In the academic domain, automatic paper-reviewer matching has become an increasingly crucial task recently due to explosive growth in the number of submissions to conferences and journals. Given a huge volume of (e.g., several thousand) submissions, it is prohibitively time-consuming for chairs or editors to manually assign papers to appropriate reviewers. Even if reviewers can self-report their expertise on certain papers through a bidding process, they can hardly scan all submissions, hence an accurate pre-ranking result should be delivered to them so that they just need to check a shortlist of papers. In other words, a precise scoring system that can automatically judge the expertise relevance between each paper and each reviewer becomes an increasingly urgent need for finding qualified reviewers.
@@
Paper-reviewer matching has been extensively studied as a text mining task (Mimno and McCallum, 2007; Charlin and Zemel, 2013; Jin et al., 2017; Anjum et al., 2019; Singh et al., 2023; Stelmakh et al., 2023), which aims to estimate to what extent a reviewer is qualified to review a submission given the text (e.g., title and abstract) of the submission as well as the papers previously written by the reviewer. Intuitively, as shown in Figure 1, there are three major factors considered by related studies. (1) Semantic: Taking the submission $p$ as a query, if the papers most semantically relevant to the query are written by a reviewer $r$, then $r$ should be qualified to review $p$. This intuition is used by previous methods such as the Toronto Paper Matching System (TPMS) (Charlin and Zemel, 2013), where tf–idf is used for calculating the semantic relevance. (2) Topic: If a reviewer $r$’s previous papers share many fine-grained research topics with the submission $p$, then $r$ is assumed to be an expert reviewer of $p$. This assumption is utilized by topic modeling approaches (Mimno and McCallum, 2007; Jin et al., 2017; Anjum et al., 2019). (3) Citation: Authors of the papers cited by the submission $p$ are more likely to be expert reviewers of $p$. This intuition is leveraged by studies (Singh et al., 2023; Stelmakh et al., 2023) using citation-enhanced scientific pre-trained language models (PLMs) such as SPECTER (Cohan et al., 2020) and SciNCL (Ostendorff et al., 2022). Note that most previous studies do not assume that the topics and references of each paper are provided as input. Instead, such information should be inferred from the paper text. [^1] The reasons why related studies make such an assumption are multifold in our view. To be specific, topics selected by the authors when they submit the paper are too coarse (e.g., “Text Mining”), while paper-reviewer matching relies heavily on more fine-grained topics (e.g., “Community Question Answering”); references in the submission do not necessarily cover all papers that ought to be cited, so we should infer what the submission should cite rather than what it actually cites.
@@
We conduct experiments on four datasets covering different fields including machine learning, computer vision, information retrieval, and data mining. Three of the datasets are released in previous studies (Mimno and McCallum, 2007; Singh et al., 2023; Karimzadehgan et al., 2008). The fourth is newly annotated by us, which is larger than the previous three and contains more recent papers. Experimental results show that our proposed CoF model consistently outperforms state-of-the-art paper-reviewer matching approaches and scientific PLM baselines on all four datasets. Further ablation studies validate the reasons why CoF is effective: (1) CoF jointly considers three factors rather than just one, (2) CoF chains the three factors to enable a progressive selection process of relevant reviewers instead of merging all factors in one step, and (3) CoF improves upon the baselines in each factor empirically.
@@ 2.1. Problem Definition
Given a set of paper submissions $\mathcal{P}=\{p_{1},p_{2},...,p_{M}\}$ and a set of candidate reviewers =$\mathcal{R}=\{r_{1},r_{2},...,r_{N}\}$, the paper-reviewer match- ing task aims to learn a function $f:\mathcal{P}\times\mathcal{R}\rightarrow\mathbb{R}$, where $f(p,r)$ reflects the expertise relevance between the paper $p$ and the reviewer $r$ (i.e., how knowledgeable that $r$ is to review $p$). We conform to the following three key assumptions made by previous studies (Mimno and McCallum, 2007; Karimzadehgan et al., 2008; Liu et al., 2014; Anjum et al., 2019; Singh et al., 2023): (1) We do not know any $f(p,r)\ (p\in\mathcal{P},r\in\mathcal{R})$ as supervision, which is a natural assumption for a fully automated paper-reviewer matching system. In other words, $f$ should be derived in a zero-shot setting, possibly by learning from available data from other sources. (2) To characterize each paper $p\in\mathcal{P}$, its text information (e.g., title and abstract) is available, denoted by $\textsf{Text}(p)$. (3) To characterize each reviewer $r\in\mathcal{R}$, its previous papers are given, denoted by $\mathcal{Q}_{r}=\{q_{r,1},q_{r,2},...,q_{r,|\mathcal{Q}_{r}|}\}$. The text information of each previous paper $q\in\mathcal{Q}_{r}$ is also provided. $\mathcal{Q}_{r}$ is called the publication profile of $r$ (Mysore et al., 2023). In practice, $\mathcal{Q}_{r}$ may be a subset of $r$’s previous papers (e.g., those published within the last 10 years or those published in top-tier venues only). To summarize, the task is defined as follows:
@@ 2.2. Semantic, Topic, and Citation Factors
Citation Factor. Recent studies (Singh et al., 2023; Stelmakh et al., 2023) adopt scientific PLMs, such as SPECTER (Cohan et al., 2020) and SciNCL (Ostendorff et al., 2022), for paper-reviewer matching. During their pre-training process, both SPECTER and SciNCL are initialized from SciBERT (Beltagy et al., 2019) and trained on a large number of citation links between papers. Empirical results show that emphasizing such citation information significantly boosts their performance in comparison with SciBERT. The motivation of considering the citation factor in paper-reviewer matching is also clear: if a paper $p$ cites many papers written by a reviewer $r$, then $r$ is more likely to be a qualified reviewer of $p$.
@@ 3.1. Chain-of-Factors Matching
We then rank all papers in $\mathcal{Q}_{\mathbb{S}}$ according to $f_{\rm topic}(p,\cdot)$ and pick those top-ranked ones as the output of this step, which we denote as $\mathcal{Q}_{\mathbb{S}\rightarrow\mathbb{T}}$, where $\mathbb{T}$ stands for the topic factor.
After checking the topic factor, we further consider citation signals. Given that two papers share common fine-grained research topics, the citation link should provide an even stronger signal of the relevance between two papers. For instance, if two papers are both about “Information Extraction”, then one citing the other may further imply that they are studying the same task or using the same dataset. However, without the premise that two papers have common research fields, the citation link becomes a weaker indicator. For example, a paper about “Information Extraction” can cite a paper about “Large Language Models” simply because the former paper uses the large language model released in the latter one. This highlights our motivation to chain the three factors for a step-by-step, coarse-to-fine selection process of relevant papers and expert reviewers. Formally, given $\mathcal{Q}_{\mathbb{S}\rightarrow\mathbb{T}}$, we use a citation-related instruction to calculate the citation-aware relevance between $p$ and each selected paper.
@@
Here, $f(p,r)$ is the final relevance score between $p$ and $r$, which can be used to rank all candidate reviewers for $p$. Note that in the last step, we consider the sum of three types of relevance, so our chain-of-factors matching strategy can be denoted as $\mathbb{S}\rightarrow\mathbb{T}\rightarrow\mathbb{S}+\mathbb{T}+\mathbb{C}$. In our experiments (Section 4.3), we will demonstrate its advantage over only considering the citation factor in the last step (i.e., $\mathbb{S}\rightarrow\mathbb{T}\rightarrow\mathbb{C}$) or simply merging all factors in one step (i.e., $\mathbb{S}+\mathbb{T}+\mathbb{C}$).
@@ 3.3. Model Training
In this section, we introduce the data and objective used to pre-trained the instruction-guided paper encoder $g(\cdot|\cdot)$.
Pre-training Data. For the semantic factor, each submission $p$ is treated as a “query” and each paper $q$ in a reviewer’s publication profile is viewed as a “document”. Thus, our $g(\cdot|{\rm semantic})$ is learned to maximize the inner product of an ad-hoc query and its semantically relevant document in the vector space. To facilitate this, we adopt the Search dataset from the SciRepEval benchmark (Singh et al., 2023) to pre-train our model, where the queries are collected from an academic search engine, and the relevant documents are derived from large-scale user click-through data.
Our $g(\cdot|{\rm topic})$ is trained to maximize the inner product of two papers $p$ and $q$ if they share common research fields. We utilize the MAPLE benchmark (Zhang et al., 2023b) as pre-training data in which millions of scientific papers are tagged with their fine-grained fields-of-study from the Microsoft Academic Graph (Sinha et al., 2015). For example, for CS papers in MAPLE, there are over 15K fine-grained research fields, and each paper is tagged with about 6 fields on average. Such data are used to derive topically relevant paper pairs.
Our $g(\cdot|{\rm citation})$ is learned to maximize the inner product of two papers $p$ and $q$ if $p$ cites $q$. Following (Cohan et al., 2020; Ostendorff et al., 2022), we leverage a large collection of citation triplets $(p,q^{+},q^{-})$ constructed by Cohan et al. (Cohan et al., 2020), where $p$ cites $q^{+}$ but does not cite $q^{-}$.
@@
Pre-training Objective. For all three factors, each sample from their pre-training data can be denoted as $(p,q^{+},q^{-}_{1},q^{-}_{2},...,q^{-}_{T})$, where $q^{+}$ is relevant to $p$ (i.e., when $\phi={\rm semantic}$, $q^{+}$ is clicked by users in a search engine given the search query $p$; when $\phi={\rm topic}$, $q^{+}$ shares fine-grained research fields with $p$; when $\phi={\rm citation}$, $q^{+}$ is cited by $p$) and $q^{-}_{t}\ (t=1,2,...,T)$ are irrelevant to $p$. Given a factor $\phi$ and its training sample $(p,q^{+},q^{-}_{1},q^{-}_{2},...,q^{-}_{T})$, we can obtain $g(p|\phi)$, $g(q^{+}|\phi)$, $g(q^{-}_{1}|\phi)$, …, and $g(q^{-}_{T}|\phi)$ using the instruction encoder ${\rm Enc}_{i}(\cdot)$ and the paper encoder ${\rm Enc}_{p}(\cdot)$. Then, we adopt a contrastive loss (Oord et al., 2018) to train our model:
@@ 4.1.1. Evaluation Datasets
  • NIPS (Mimno and McCallum, 2007) is a pioneering benchmark dataset for paper-reviewer matching. It consists of expertise relevance judgements between 34 papers accepted by NIPS 2006 and 190 reviewers. Annotations were done by 9 researchers from the NIPS community, and the score of each annotated paper-reviewer pair can be “3” (very relevant), “2” (relevant), “1” (slightly relevant), or “0” (irrelevant). Note that for each paper, the annotators only judge its relevance with a subset of reviewers.
  • SciRepEval (Singh et al., 2023) is a comprehensive benchmark for evaluating scientific document representation learning methods. Its paper-reviewer matching dataset combines the annotation effort from multiple sources (Mimno and McCallum, 2007; Liu et al., 2014; Zhao et al., 2022). Specifically, Liu et al. (Liu et al., 2014) added relevance scores of more paper-reviewer pairs to the NIPS dataset to mitigate its annotation sparsity; Zhao et al. (Zhao et al., 2022) provided some paper-reviewer relevance ratings for the ICIP 2016 conference. The combined dataset still adopts the “0”-“3” rating scale.
  • SIGIR (Karimzadehgan et al., 2008) contains 73 papers accepted by SIGIR 2007 and 189 prospective reviewers. Instead of annotating each specific paper-reviewer pair, the dataset constructors assign one or more aspects of information retrieval (e.g., “Evaluation”, “Web IR”, and “Language Models”, with 25 candidate aspects in total) to each paper and each reviewer. Then, the relevance between a paper and a reviewer is determined by their aspect-level similarity. In our experiments, to align with the rating scale in NIPS and SciRepEval, we discretize the Jaccard similarity between a paper’s aspects and a reviewer’s aspects to map their relevance to “0”-“3”.